OpenStudy (anonymous):

The derivative of the 3rd sqrt of x = ? The derivative of the 3rd sqrt of x = ? @Mathematics

OpenStudy (amistre64):

derivative of x^(1/3), and apply the power rule

OpenStudy (anonymous):

as in \[\sqrt[3]{x}\]

OpenStudy (amistre64):

radicals are fractional exponents

OpenStudy (anonymous):

First understand that the cube root of a number is the same as that number raised to 1/3. This makes it easy. Multiply by 1/3 and subtract 1 from 1/3 for the new exponent. So you'd get 1/3x^-2/3

OpenStudy (amistre64):

unless your asking for the limit definition of a derivative

OpenStudy (amistre64):

then its messier :)

OpenStudy (anonymous):

as in \[\sqrt[3]{x}\]

OpenStudy (anonymous):

ok and rewritten, would that be \[1\div(3\sqrt[3]{x ^{2}})\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well actually x shouldn't be in the denominator I don't think.

OpenStudy (amistre64):

x goes negative on an exponent, its denomed

OpenStudy (anonymous):

oops i'm stupid

OpenStudy (anonymous):

so, what is the answer then?

OpenStudy (anonymous):

What you had, you were 100% correct!

OpenStudy (anonymous):

ah ok, tanks again :)